Find out median for the following data

x: 10,15,18,23,27,30,31 f

20,30,40,50,60,70



1
Expert's answer
2022-03-29T08:46:43-0400

"Sorted\,\,sample:\\10,15,18,20,23,27,30,30,31,40,50,60,70\\There\,\,are\,\,13 elements\,\,\left( odd\,\,number \right) , hence\\Med={x^*}_{\frac{13+1}{2}}={x^*}_7=30"